OpenStudy (anonymous):

Which functions are exponential decay functions? Choose exactly two answers that are correct.

OpenStudy (anonymous):

f(x)=0.9^x g(x)=(6/5)^x h(x)=1.25^x j(x)=(4/5)^x

OpenStudy (anonymous):

Can you tell me what exponential decay is?

OpenStudy (anonymous):

No. I mean I guess I could look it up in the dictionary, but at the top of my head; no.

OpenStudy (anonymous):

Okay then let's think about this from another approach, do you know what an exponential function is? If something increases exponentially, what happens to y as x increases by a certain amount each time.

OpenStudy (anonymous):

OK...

OpenStudy (anonymous):

The second sentence was meant to be a question (missed off the question mark though, sorry!).

OpenStudy (anonymous):

That is a question that I cannot answer. ;)

OpenStudy (anonymous):

lol okay, sounds like someone wasn't paying attention in their maths lesson! We're looking for a function that vaguely resembles this Now, let's have a think about your functions: f(x)=0.9^x f(1)=0.9 f(2)=0.81 f(3)=0.729 g(x)=(6/5)^x g(1)=1.2 g(2)=1.44 g(3)=1.728 h(x)=1.25^x h(1)=1.25 h(2)=1.5625 h(3)=1.953125 j(x)=(4/5)^x j(1)=0.8 j(2)=0.64 j(3)=0.512 What can you see about these functions as x increases?

OpenStudy (anonymous):

@tom982 Hold on, It might take me a minute to figure this out. Also technically I didn't even go to my math class (It's online) I'm in the process of moving and my computer had to be shipped back to my school and I have a ton of other issues going on. :)

OpenStudy (anonymous):

I think it's D.

OpenStudy (anonymous):

I understand, well if you get stuck with any of your other things feel free to shoot me a message and I'll do my best to explain. D is correct, but why? What's happening with j(x) as x increases?

OpenStudy (anonymous):

There's suppose to be one more but I don't know which. @tom982

OpenStudy (anonymous):

I know there is, we're getting close to that. Can you answer my previous questions?

OpenStudy (anonymous):